OpenStudy (anonymous):

What is the equation of the line that passes through the points (−2, 1) and (1, 10)? 3x − y = −7 3x − y = −5 3x − y = 5 x + 3y = −5

OpenStudy (anonymous):

plug those points into each equation and see which one holds true.

OpenStudy (anonymous):

 you coud graph it, or get the slope -> 10-1/1--2 (or 1+2) and get 9/3 which is 3 so the slope is 3. Plug in the #s to the point slope equation--- y-1=3(x+2) and simplify to y-1=3x+6 and get y=3x+5 and put it into standard form-- -3x+y=5 and then multiply everything by -1 to get 3x-y=-5, so the second answer is correct! TAH DAH, MATH